Question 3.

The Timber Ridge Company has the following relationships:

Sales/Total assets = 3.80; ROA = 0.0980

What is Timber Ridge’s net profit margin? Round to 3 decimal places.

Question 5

Archware Systems has total assets of $35.594
billion, total debt of $9.678 billion, and net sales of $23.720 billion. Their
net profit margin for the year was 0.22, while the operating profit
margin was 30 percent. What is Archware’s net income? (Answer needs to be stated in billions. For example: 2.83) Round to two decimal places.

QUESTION 6

Which one of the following statements is NOT true?

Question 6 options:

DSO measures in days, the time the firm takes to convert its receivables into cash.

One ratio that measures the efficiency of a firm's collection policy is days' sales outstanding.

The accounts receivables turnover ratio measures how quickly the firm collects on its credit sales.

The more days that it takes the firm to collect on its receivables, the more efficient the firm is.

QUESTION 7

Blue Air Inc., has net sales of $791,000 and accounts receivables of $169,000. What are the firm's accounts receivables turnover? Round to two decimal places.

QUESTION 8

JP Vineyards has sales of $820,000, a gross profit margin of 0.330, and inventory of $171,000. What is the company's inventory turnover ratio? Round to two decimal places.
Answer #1

3.

Profit margin = roa/(sales/total assets) =0.098/3.98=0.025

please ask remaining parts seperately